# Unsupervised Learning for Quadratic Assignment

### v1

@@ 1 Introduction
Combinatorial Optimization (CO) represents a central challenge in computer science and operations research, targeting optimal solutions within a large search space. CO encompasses numerous practical applications, including transportation logistics, production scheduling, network design, and resource allocation. The computational complexity of CO problems is typically NP-hard, making exact methods intractable for large instances. Researchers have thus developed approximation algorithms, heuristics, and hybrid approaches that balance solution quality with computational feasibility. These methods include simulated annealing, genetic algorithms, tabu search, and various problem-specific heuristics that can generate high-quality solutions within reasonable time budget Kirkpatrick et al. (1983); Holland (1992); Johnson and McGeoch (1997); Glover and Laguna (1998); Gomes and Selman (2001); Blum and Roli (2003).
@@ 1.1 Data-driven Combinatorial Optimization
Recently, data-driven methods have gained significant attention in addressing combinatorial optimization problems. Taking the Travelling Salesman Problem (TSP) as an example, researchers have explored both Supervised Learning (SL) and Reinforcement Learning (RL) methods. In SL, approaches such as pointer networks and graph neural networks attempt to learn mappings from problem instances to solutions by training on optimal or near-optimal tours Joshi et al. (2019); Vinyals et al. (2015). These models learn to imitate optimal or near-optimal solutions, leading to significant computational expense when building the training dataset. RL approaches frame TSP as a sequential decision-making problem where an agent learns to construct tours in a Markov decision process framework. While RL models have shown promise in small instances, these methods face significant challenges when scaling to larger problems. Furthermore, the sparse rewards and the high variance during training make it difficult for RL agents to learn effective policies Bello et al. (2016).
An alternative approach is Unsupervised Learning (UL), which avoids sequential decision making and does not require labelled data. In  Min et al. (2023), the authors propose a surrogate loss for the TSP objective by using a soft indicator matrix $\mathbb{T}$ to construct the Hamiltonian cycle, which is then optimized for minimum total distance. The $\mathbb{T}$ operator can be essentially interpreted as a soft permutation operator, as demonstrated in Min and Gomes (2023), which represents a rearrangement of nodes on the route $1\rightarrow 2\rightarrow 3\rightarrow\cdots\rightarrow n\rightarrow 1$, with $n$ representing the number of cities.
@@
$\mathcal{L}_{\text{TSP}}=\langle\mathbb{T}\mathbb{V}\mathbb{T}^{T},\mathbf{D}_{\text{TSP}}\rangle,$ (1)
where $\mathbb{V}$ represents a Hamiltonian cycle matrix encoding the route $1\rightarrow 2\rightarrow\cdots\rightarrow n\rightarrow 1$, $\mathbb{T}$ is an approximation of a hard permutation matrix $\mathbf{P},$ and $\mathbf{D}_{\text{TSP}}$ is the distance matrix with self-loop distances set to $\lambda$. $\mathbb{T}\mathbb{V}\mathbb{T}^{T}$ is a heat map that represents the probability that each edge belongs to the optimal solution, which guides the subsequent search process.
@@ 1.2 Quadratic Assignment Problem
The Quadratic Assignment Problem (QAP), essentially a permutation optimization problem, is an NP-hard problem with numerous applications across facility layout, scheduling, and computing systems. For example, in facility layout, QAP finds the optimal permutation of facilities minimizing total interaction costs based on inter-facility flows and inter-location distances. QAP’s applications also include manufacturing plant design, healthcare facility planning, VLSI circuit design, telecommunications network optimization, and resource scheduling Koopmans and Beckmann (1957); Lawler (1963).
@@
$\min_{\mathbf{P}\in S_{n}}\langle\mathbf{P}\mathbf{F}\mathbf{P}^{T},\mathbf{D}\rangle,$ (2)
@@ 1.3 Unsupervised Learning for QAP
In this paper, we use PLUME search to solve QAP. Following the TSP fashion in Min et al. (2023), we use neural networks (NNs) coupled with a Gumbel-Sinkhorn operator to construct a soft permutation matrix $\mathbb{T}$ that approximates a hard permutation matrix $\mathbf{P}$. We use the soft permutation $\mathbb{T}$ to guide the subsequent search process, as shown in Figure 1.
@@
Here, for QAP, while TSP uses the heat map $\mathcal{H}=\mathbb{T}\mathbb{V}\mathbb{T}^{T}$ as a soft matrix (heat map) to guide the search, we directly decode a hard permutation matrix $\mathbf{P}$ from $\mathbb{T}$. This permutation matrix $\mathbf{P}$ represents an alignment in QAP and serves as an initialization to guide the subsequent search.
@@ 2 Tabu Search
We adopt Tabu search as the backbone of our PLUME search framework. Tabu Search (TS) is a method introduced by Glover Glover and Laguna (1998) that enhances local search methods by employing memory structures to navigate the solution space effectively. Unlike traditional hill-climbing algorithms, TS allows non-improving moves to escape local optima and uses adaptive memory to avoid cycling through previously visited solutions. The algorithm maintains a tabu list that prohibits certain moves, creating a dynamic balance between intensification and diversification strategies.
@@
TS has emerged as one of the most effective metaheuristics for addressing QAP instances Taillard (1991); Drezner (2003); James et al. (2009). The algorithm navigates the solution space through strategic move evaluations and maintaining memory structures to prevent cycling and encourage diversification. For QAP, TS typically begins with a random permutation as the initial solution and employs a swap-based neighborhood structure where adjacent solutions are generated by exchanging the assignments of two facilities. A key component of TS is the tabu list, which records recently visited solutions or solution attributes to prevent immediate revisiting. In QAP implementations, the tabu list typically tracks recently swapped facility pairs, prohibiting their re-exchange for a specified number of iterations—the tabu tenure. This memory structure forces the search to explore new regions of the solution space even when immediate improvements are not available, helping the algorithm escape local optima. The performance depends mainly on three parameters: neighbourhoodSize, which controls the sampling density from the complete swap neighbourhood at each iteration; evaluations, which establishes the maximum computational budget as measured by objective function calculations; and maxFails, which implements an adaptive early termination criterion that halts the search after a predefined number of consecutive non-improving iterations. Together, these parameters balance exploration breadth against computational efficiency, ensuring both effective solution space coverage and predictable runtime performance Glover (1989); Blum and Roli (2003); Battiti and Tecchiolli (1994); Misevicius (2005).
@@ 3 Model
We design NNs that preserve permutation equivalence while effectively encoding the underlying structure of QAP. Specifically, our NNs preserve permutation equivariance to ensure consistent handling of isomorphic problem instances. Furthermore, our model generalizes across variable problem sizes without modifying the NNs.
3.1 Permutation-Equivariant QAP Architecture
The NN’s structure is shown in Figure 2. Our permutation-equivariant architecture has two inputs: node positions $\mathbf{X}\in\mathbb{R}^{n\times 2}$ representing 2D coordinates and flow matrices $\mathbf{F}\in\mathbb{R}^{n\times n}$ representing pairwise interactions, where $n$ is the number of facilities. Our model consists of three main components: a position encoder, two flow feature encoders, and the main network. The position encoder maps facility coordinates through a shared Multi-Layer Perceptron (MLP) $\phi_{\text{pos}}$ to embed spatial information, the flow encoders process row-wise and column-wise interactions through pooling operations and MLPs:
$\mathbf{H}_{\text{pos}}=\phi_{\text{pos}}(\mathbf{X}),\quad\mathbf{H}_{\text{row}}=\psi_{\text{row}}\left(\sum_{j=1}^{N}\phi_{\text{row}}(\mathbf{F}_{ij})\right),\quad\mathbf{H}_{\text{col}}=\psi_{\text{col}}\left(\sum_{i=1}^{N}\phi_{\text{col}}(\mathbf{F}_{ij})\right),$ (3)
where $\phi_{\text{row}},\phi_{\text{col}}:\mathbb{R}\rightarrow\mathbb{R}^{d}$ and $\phi_{\text{pos}}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{d}$ are MLPs that transform input features into $d$-dimensional latent representations, and $\psi_{\text{row}},\psi_{\text{col}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ are MLPs that process the aggregated features to capture facility interactions. These features are concatenated and processed through $L$ permutation-equivariant MLP layers before the final projection:
$\mathbf{H}^{(0)}=[\mathbf{H}_{\text{pos}}\|\mathbf{H}_{\text{row}}\|\mathbf{H}_{\text{col}}],\quad\mathbf{H}^{(\ell+1)}=\text{MLP}^{(\ell)}(\mathbf{H}^{(\ell)}),\quad\mathbf{Y}=\Theta\mathbf{H}^{(L)},$ (4)
where $\mathbf{H}^{(0)}\in\mathbb{R}^{n\times 3d}$ contains the concatenated initial node features, $\mathbf{H}^{(\ell)}$ represents node embeddings at layer $\ell$, and $\text{MLP}^{(\ell)}$ denotes the $\ell$-th layer. These MLPs map from $3d$ to $d$ dimensions, while the final linear layer $\Theta\in\mathbb{R}^{d\times m}$ projects from $d$ to $m$ dimensions. $\mathbf{Y}\in\mathbb{R}^{n\times m}$ is the final output logits for assignment prediction. This architecture preserves permutation equivariance, that is, $f(\Pi\mathbf{X},\Pi\mathbf{F}\Pi^{T})=\Pi f(\mathbf{X},\mathbf{F})$ for any permutation matrix $\Pi\in S_{n}$, where $f$ represents the network function mapping inputs $\mathbf{X},\mathbf{F}$ to output $\mathbf{Y}$.
@@ 3.2 Building Soft Permutation $\mathbb{T}$
$\mathcal{F}=\alpha\tanh(\mathbf{Y}\mathbf{Y}^{T}),$ (5)
@@
During inference, we obtain a discrete permutation matrix $\mathbf{P}$ by applying the Hungarian algorithm to the scaled logits: $\mathbf{P}=\text{Hungarian}(-\frac{\mathcal{F}+\gamma\times\text{Gumbel noise}}{\tau})$. We use the CUDA implementation of the batched linear assignment solver for the Hungarian operator from  Karpukhin et al. (2024).
@@ 3.3 Invariance Property of Permutation Representation
The soft permutation matrix $\mathbb{T}$ is constructed through Equations 5 and 6 to preserve permutation equivariance. Let $\Pi\in S_{n}$ represent a random permutation on the nodes, the distance and flow matrices transform as $\mathbf{D}=\Pi\mathbf{D}_{0}\Pi^{T}$ and $\mathbf{F}=\Pi\mathbf{F}_{0}\Pi^{T}$, where $\mathbf{D}_{0}$ and $\mathbf{F}_{0}$ are the distance matrix and flow matrix before this random permutation respectively. Now, let $\mathbf{Y}_{0}$ denote the initial output, given that the network output transforms under permutation $\Pi$ as $\mathbf{Y}=\Pi\mathbf{Y}_{0}$, we then have $\mathbb{T}=\Pi\mathbb{T}_{0}\Pi^{T}$. Consequently, the objective function $\langle\mathbb{T}\mathbf{F}\mathbb{T}^{T},\mathbf{D}\rangle$ remains invariant. To be specific, $\forall~{}\Pi\in S_{n}$, we have:
$\langle\mathbb{T}\mathbf{F}\mathbb{T}^{T},\mathbf{D}\rangle$ $=\langle\Pi\mathbb{T}_{0}\Pi^{T}\Bigl{(}\Pi\mathbf{F}_{0}\Pi^{T}\Bigr{)}\Bigl{(}\Pi\mathbb{T}_{0}\Pi^{T}\Bigr{)}^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle$ (7)
$=\langle\Pi\mathbb{T}_{0}\underbrace{\Pi^{T}\Pi}_{=I}\mathbf{F}_{0}\underbrace{\Pi^{T}\Pi}_{=I}\mathbb{T}^{T}_{0}\Pi^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle=\langle\Pi\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{T}_{0}\Pi^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle$
$=\langle\Pi\Bigl{(}\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{T}_{0}\Bigr{)}\Pi^{T},\Pi\mathbf{D}_{0}\Pi^{T}\rangle=\langle\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{T}_{0},\mathbf{D}_{0}\rangle.$
This invariance property guarantees consistent solutions for isomorphic problem instances. Overall, Equations 5 and 6 preserve permutation symmetry while enabling gradient-based optimization, with the additional benefit of allowing the model to naturally generalize across different problem sizes. This generalization capability arises because the soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$ always matches the input graph size, independent of the model’s parameters. Consequently, our approach scales to problems of varying sizes without needing to modify the NN.

### v2

@@ 1 Introduction
Combinatorial Optimization (CO) represents a central challenge in computer science and operations research, targeting optimal solutions within a large search space. CO encompasses numerous practical applications, including transportation logistics, production scheduling, network design, and resource allocation. The computational complexity of CO problems is typically NP-hard, making exact methods intractable for large instances. Researchers have thus developed approximation algorithms, heuristics, and hybrid approaches that balance solution quality with computational feasibility. These methods include simulated annealing, genetic algorithms, tabu search, and various problem-specific heuristics that can generate high-quality solutions within reasonable time budget (Kirkpatrick et al., 1983; Holland, 1992; Johnson and McGeoch, 1997; Glover and Laguna, 1998; Gomes and Selman, 2001; Blum and Roli, 2003).
@@ 1.1 Data-driven Combinatorial Optimization
Recently, data-driven methods have gained significant attention in addressing combinatorial optimization problems. Taking the Travelling Salesman Problem (TSP) as an example, researchers have explored both Supervised Learning (SL) and Reinforcement Learning (RL) methods. In SL, approaches such as pointer networks and graph neural networks attempt to learn mappings from problem instances to solutions by training on optimal or near-optimal tours (Joshi et al., 2019; Vinyals et al., 2015). These models learn to imitate optimal or near-optimal solutions, leading to significant computational expense when building the training dataset. RL approaches frame TSP as a sequential decision-making problem where an agent learns to construct tours in a Markov decision process framework. While RL models have shown promise in small instances, these methods face significant challenges when scaling to larger problems. Furthermore, the sparse rewards and the high variance during training make it difficult for RL agents to learn effective policies (Bello et al., 2016).
An alternative approach is Unsupervised Learning (UL), which avoids sequential decision making and does not require labelled data. In  (Min et al., 2023), the authors propose a surrogate loss for the TSP objective by using a soft indicator matrix $\mathbb{T}$ to construct the Hamiltonian cycle, which is then optimized for minimum total distance. The $\mathbb{T}$ operator can be essentially interpreted as a soft permutation operator, as demonstrated in (Min and Gomes, 2023), which represents a rearrangement of nodes on the route $1\rightarrow 2\rightarrow 3\rightarrow\cdots\rightarrow n\rightarrow 1$, with $n$ representing the number of cities.
@@
$\mathcal{L}_{\text{TSP}}=\langle\mathbb{T}\mathbb{V}\mathbb{T}^{\top},\mathbf{D}_{\text{TSP}}\rangle,$ (1)
where $\mathbb{V}$ represents a Hamiltonian cycle matrix encoding the route $1\rightarrow 2\rightarrow\cdots\rightarrow n\rightarrow 1$, $\mathbb{T}$ is an approximation of a hard permutation matrix $\mathbf{P},$ and $\mathbf{D}_{\text{TSP}}$ is the distance matrix with self-loop distances set to $\lambda$. $\mathbb{T}\mathbb{V}\mathbb{T}^{\top}$ is a heat map that represents the probability that each edge belongs to the optimal solution, which guides the subsequent search process.
@@ 1.2 Quadratic Assignment Problem
The Quadratic Assignment Problem (QAP), essentially a permutation optimization problem, is an NP-hard problem with numerous applications across facility layout, scheduling, and computing systems. For example, in facility layout, QAP finds the optimal permutation of facilities minimizing total interaction costs based on inter-facility flows and inter-location distances. QAP’s applications also include manufacturing plant design, healthcare facility planning, VLSI circuit design, telecommunications network optimization, and resource scheduling (Koopmans and Beckmann, 1957; Lawler, 1963).
@@
$\min_{\mathbf{P}\in S_{n}}\langle\mathbf{P}\mathbf{F}\mathbf{P}^{\top},\mathbf{D}\rangle,$ (2)
@@ 1.3 Unsupervised Learning for QAP
In this paper, we use PLUME search to solve QAP. Following the TSP fashion in (Min et al., 2023), we use neural networks (NNs) coupled with a Gumbel-Sinkhorn operator to construct a soft permutation matrix $\mathbb{T}$ that approximates a hard permutation matrix $\mathbf{P}$. We use the soft permutation $\mathbb{T}$ to guide the subsequent search process, as shown in Figure 1.
@@
Here, for QAP, while TSP uses the heat map $\mathcal{H}=\mathbb{T}\mathbb{V}\mathbb{T}^{\top}$ as a soft matrix (heat map) to guide the search, we directly decode a hard permutation matrix $\mathbf{P}$ from $\mathbb{T}$. This permutation matrix $\mathbf{P}$ represents an alignment in QAP and serves as an initialization to guide the subsequent search.
@@ 2 Tabu Search
We adopt Tabu search as the backbone of our PLUME search framework. Tabu Search (TS) is a method introduced by Glover (Glover and Laguna, 1998) that enhances local search methods by employing memory structures to navigate the solution space effectively. Unlike traditional hill-climbing algorithms, TS allows non-improving moves to escape local optima and uses adaptive memory to avoid cycling through previously visited solutions. The algorithm maintains a tabu list that prohibits certain moves, creating a dynamic balance between intensification and diversification strategies.
@@
TS has emerged as one of the most effective metaheuristics for addressing QAP instances (Taillard, 1991; Drezner, 2003; James et al., 2009). The algorithm navigates the solution space through strategic move evaluations and maintaining memory structures to prevent cycling and encourage diversification. For QAP, TS typically begins with a random permutation as the initial solution and employs a swap-based neighborhood structure where adjacent solutions are generated by exchanging the assignments of two facilities. A key component of TS is the tabu list, which records recently visited solutions or solution attributes to prevent immediate revisiting. In QAP implementations, the tabu list typically tracks recently swapped facility pairs, prohibiting their re-exchange for a specified number of iterations—the tabu tenure. This memory structure forces the search to explore new regions of the solution space even when immediate improvements are not available, helping the algorithm escape local optima. The performance depends mainly on three parameters: neighbourhoodSize, which controls the sampling density from the complete swap neighbourhood at each iteration; evaluations, which establishes the maximum computational budget as measured by objective function calculations; and maxFails, which implements an adaptive early termination criterion that halts the search after a predefined number of consecutive non-improving iterations. Together, these parameters balance exploration breadth against computational efficiency, ensuring both effective solution space coverage and predictable runtime performance (Glover, 1989; Blum and Roli, 2003; Battiti and Tecchiolli, 1994; Misevicius, 2005).
@@ 3 Model
We propose a permutation-equivariant neural architecture for the QAP. The network jointly encodes (i) pairwise flows among facilities and (ii) pairwise distances among locations, and fuses these with per-node positional features. Permutation equivariance is enforced by construction through row-wise symmetric pooling operators (sum, mean, max), shared multilayer perceptron (MLP) encoders for facility–facility and location–location interactions, and equivariant message passing updates that avoid in-place modifications to ensure stable gradient flow.
@@
As shown in Figure 2, our neural network consists of a facility encoder and a location encoder, each based on lightweight MLPs and pooled symmetric operators, a positional lifting module that embeds 22D coordinates into the hidden space, and a fusion block that integrates facility, location, and positional features across multiple layers. By construction, our model guarantees permutation equivariance, in the sense that permuting rows and columns of the inputs permutes the learned embeddings in the same way.
Let $n$ denote the number of facilities/locations. The inputs of our model are 2D coordinates $\mathbf{X}\in\mathbb{R}^{n\times 2}$ and a (typically symmetric) flow matrix $\mathbf{F}\in\mathbb{R}^{n\times n}$. We precompute once per batch the Euclidean distance matrix $\mathbf{D}\in\mathbb{R}^{n\times n}.$ $\mathbf{X}$ is embedded once by a 2-layer MLP $\phi_{x}:\mathbb{R}^{2}\!\to\!\mathbb{R}^{d}$,
$\mathbf{H}^{\text{pos}}=\phi_{x}(\mathbf{X})\in\mathbb{R}^{n\times d},$ (3)
and reused across layers as fixed positional context.
3.1 Encoders
The encoders provide node-aligned embeddings of flows, distances, and coordinates that respect permutation equivariance. Each encoder lifts scalar pairwise inputs into $\mathbb{R}^{d}$, aggregates across neighbors using symmetric pooling (sum/mean/max), mixes the pooled summary with an MLP, and applies a row-stochastic message pass. This design ensures that both global context and local interactions are captured before embeddings are fused downstream.
Our flow encoder transforms entries of the flow matrix $\mathbf{F}\in\mathbb{R}^{n\times n}$ with a shared 2-layer MLP $\phi_{f}:\mathbb{R}\!\to\!\mathbb{R}^{d}$: $\mathcal{X}_{ij}=\phi_{f}(\mathbf{F}_{ij})$. For each node $i$, pooled statistics over its neighbors are computed:
$\mathbf{s}^{\Sigma}_{i}=\sum_{j}\mathcal{X}_{ij},\quad\mathbf{s}^{\mu}_{i}=\tfrac{1}{n}\mathbf{s}^{\Sigma}_{i},\quad\mathbf{s}^{\max}_{i}=\max_{j}\mathcal{X}_{ij}.$ (4)
These are concatenated into $\mathbf{h}^{\text{pool}}_{i}=[\mathbf{s}^{\Sigma}_{i}\;\|\;\mathbf{s}^{\mu}_{i}\;\|\;\mathbf{s}^{\max}_{i}]\in\mathbb{R}^{3d}$ and passed through a mixing MLP $\mathrm{mix}_{f}:\mathbb{R}^{3d}\!\to\!\mathbb{R}^{d}$ to obtain $\tilde{\mathbf{h}}_{i}$. To propagate information along flow structure, we normalize $\mathbf{F}$ row-wise:
$\mathbf{W}_{F}\;=\;\mathrm{normalize}_{1}(\mathbf{F}),$ (5)
and apply a linear message map $\mathrm{msg}_{f}:\mathbb{R}^{d}\!\to\!\mathbb{R}^{d}$, giving the encoder output
$\mathbf{H}^{\text{fac}}=\tilde{\mathbf{h}}+\mathbf{W}_{F}\,\mathrm{msg}_{f}(\tilde{\mathbf{h}})\in\mathbb{R}^{n\times d}.$ (6)
Thus, each facility embedding combines local pooled features with weighted messages from other facilities, scaled by flow magnitude.
Our location encoder is structurally identical but uses distances. A shared 2-layer MLP $\phi_{\ell}:\mathbb{R}\!\to\!\mathbb{R}^{d}$ embeds distances into $\mathcal{G}_{ij}=\phi_{\ell}(D_{ij})$, producing pooled summaries $\mathbf{z}^{\text{pool}}_{i}\in\mathbb{R}^{3d}$. After mixing, we obtain $\tilde{\mathbf{z}}=\mathrm{mix}_{\ell}(\mathbf{z}^{\text{pool}})$. To emphasize nearby nodes, we construct a kernel from the inverse distance:
$\mathbf{W}_{D}=\mathrm{normalize}_{1}\!\big{(}(\mathbf{D}+\epsilon\mathbf{1})^{-1}\big{)},\qquad\epsilon{=}10^{-3},$ (7)
and update embeddings via
$\mathbf{H}^{\text{loc}}=\tilde{\mathbf{z}}+\mathbf{W}_{D}\,\mathrm{msg}_{\ell}(\tilde{\mathbf{z}})\in\mathbb{R}^{n\times d}.$ (8)
This encourages each node to aggregate more strongly from spatially close locations, consistent with the QAP cost structure.
Finally, raw coordinates are embedded by a 2-layer position lift MLP $\phi_{x}:\mathbb{R}^{2}\!\to\!\mathbb{R}^{d}$:
$\mathbf{H}^{\text{pos}}=\phi_{x}(\mathbf{X})\in\mathbb{R}^{n\times d},$ (9)
computed once and reused across layers. These positional embeddings act as fixed context that complements the dynamic flow and distance streams.
3.2 Fusion Stack and State Update
To enable effective message passing, we introduce a Fusion Stack combined with a State Update mechanism. The key idea is to iteratively refine node embeddings so that information propagates across the entire graph, thereby capturing multi-hop dependencies between facilities and locations. This is crucial for QAP, since the cost of assigning a facility to a location depends not only on its direct interaction with a single other facility but also on indirect chains of interactions involving the rest of the system.
At layer $\ell$, we construct node-aligned streams for facilities, locations, and positions, concatenate them, and fuse with a learned transformation:
$\mathbf{U}^{(\ell)}=\big{[}\mathbf{H}^{\text{fac}}\,\|\,\mathbf{H}^{\text{loc}}\,\|\,\mathbf{H}^{\text{pos}}\big{]}\in\mathbb{R}^{n\times 3d},\qquad\mathbf{H}^{(\ell)}=\mathrm{fuse}(\mathbf{U}^{(\ell)})\in\mathbb{R}^{n\times d},$ (10)
where $\mathrm{fuse}:\mathbb{R}^{3d}\!\to\!\mathbb{R}^{d}$ is implemented as a 3-layer MLP. We then perform a state update by setting
$\mathbf{H}^{\text{fac}}\leftarrow\mathbf{H}^{(\ell)},\qquad\mathbf{H}^{\text{loc}}\leftarrow\mathbf{H}^{(\ell)},$ (11)
so that both facility and location streams share the updated node representation, while $\mathbf{H}^{\text{pos}}$ remains fixed as positional context. After $n_{\text{layers}}$ layers of Fusion and State Update, the network outputs embeddings
$\mathbf{Y}=\mathbf{H}^{(n_{\text{layers}})}\in\mathbb{R}^{n\times d}.$ (12)
An additional advantage of this design is that the weighting of message updates is directly aligned with the QAP objective. We apply row normalization to the flow matrix $\mathbf{F}$ and to the inverse distance matrix $(\mathbf{D}+\epsilon)^{-1}$, yielding row-stochastic kernels that act as attention weights. This ensures that messages are propagated preferentially along cost-critical interactions: strong flows between facilities and short distances between locations. As a result, the network allocates representational capacity to the most influential dependencies while still retaining global structural context through symmetric pooling. This balance between broad structural awareness and targeted local refinement enables the model to capture higher-order dependencies that are essential for achieving globally optimal assignments.
@@ 3.3 Building Soft Permutation $\mathbb{T}$
$\mathcal{F}=\alpha\tanh(\mathbf{Y}\mathbf{Y}^{\top}),$ (13)
@@
During inference, we obtain a discrete permutation matrix $\mathbf{P}$ by applying the Hungarian algorithm to the scaled logits: $\mathbf{P}=\text{Hungarian}(-\frac{\mathcal{F}+\gamma\times\text{Gumbel noise}}{\tau})$. We use the CUDA implementation of the batched linear assignment solver for the Hungarian operator from  (Karpukhin et al., 2024).
@@ 3.4 Invariance Property of Permutation Representation
The soft permutation matrix $\mathbb{T}$ is constructed through Equations 13 and 14 to preserve permutation equivariance. Let $\Pi\in S_{n}$ represent a random permutation on the nodes, the distance and flow matrices transform as $\mathbf{D}=\Pi\mathbf{D}_{0}\Pi^{\top}$ and $\mathbf{F}=\Pi\mathbf{F}_{0}\Pi^{\top}$, where $\mathbf{D}_{0}$ and $\mathbf{F}_{0}$ are the distance matrix and flow matrix before this random permutation respectively. Now, let $\mathbf{Y}_{0}$ denote the initial output, given that the network output transforms under permutation $\Pi$ as $\mathbf{Y}=\Pi\mathbf{Y}_{0}$, we then have $\mathbb{T}=\Pi\mathbb{T}_{0}\Pi^{\top}$. Consequently, the objective function $\langle\mathbb{T}\mathbf{F}\mathbb{T}^{\top},\mathbf{D}\rangle$ remains invariant. To be specific, $\forall~\Pi\in S_{n}$, we have:
$\langle\mathbb{T}\mathbf{F}\mathbb{T}^{\top},\mathbf{D}\rangle$ $=\langle\Pi\mathbb{T}_{0}\Pi^{\top}\Bigl{(}\Pi\mathbf{F}_{0}\Pi^{\top}\Bigr{)}\Bigl{(}\Pi\mathbb{T}_{0}\Pi^{\top}\Bigr{)}^{\top},\Pi\mathbf{D}_{0}\Pi^{\top}\rangle$ (15)
$=\langle\Pi\mathbb{T}_{0}\underbrace{\Pi^{\top}\Pi}_{=I}\mathbf{F}_{0}\underbrace{\Pi^{\top}\Pi}_{=I}\mathbb{T}^{\top}_{0}\Pi^{\top},\Pi\mathbf{D}_{0}\Pi^{\top}\rangle=\langle\Pi\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{\top}_{0}\Pi^{\top},\Pi\mathbf{D}_{0}\Pi^{\top}\rangle$
$=\langle\Pi\Bigl{(}\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{\top}_{0}\Bigr{)}\Pi^{\top},\Pi\mathbf{D}_{0}\Pi^{\top}\rangle=\langle\mathbb{T}_{0}\mathbf{F}_{0}\mathbb{T}^{\top}_{0},\mathbf{D}_{0}\rangle.$
This invariance property guarantees consistent solutions for isomorphic problem instances. Overall, Equations 13 and 14 preserve permutation symmetry while enabling gradient-based optimization, with the additional benefit of allowing the model to naturally generalize across different problem sizes. This generalization capability arises because the soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$ always matches the input graph size, independent of the model’s parameters. Consequently, our approach scales to problems of varying sizes without needing to modify the NN.